.A planning machine works in two strokes namely cutting stroke and non-cutting
stroke. The total time taken for both the strokes is 2 minutes. The displacement of
the cutting stroke of the machine is given by the expression x+2y+z; whereas x is
150 cm; y is 60% lesser than x; z is average of x and y. During the cutting stroke,
initially the planning table accelerates uniformly for a displacement equal to 75%
of x from rest; then the speed remains constant for a displacement equal to 30%
of cutting stroke and retards uniformly for the remaining stroke till it comes to
stop. Find the maximum cutting velocity of the machine in cm/s, if time taken for
cutting stroke is 60% of the total time taken.
Answer #1

solve for 3 different intervals and find the final solutionGiven total time for both the strokes = 2 minutes Let displacement is d, le. d= x + 2y tz where x = 150 cm y = (2- 60 x x ) 1